OpenStudy (anonymous):

please help to find cubic equation when f' (x) =0 at(-1;-3,5) f'(x)<0 for x<-1 and x>2 f'(x)>0 for -1

OpenStudy (anonymous):

use cardano's method given below

OpenStudy (anonymous):

ok you know your derivative is negative then positive then negative, so it is a parabola facing down. the zeros are -1 and 2 so it looks like \[f'(x)=a(x+1)(x-2)\] where a is some negative number. could even be -1 who knows.

OpenStudy (amistre64):

is the derivative the cubic? it has 3 zeros...

OpenStudy (amistre64):

f' = 0 at -1,-3,5 ?

OpenStudy (anonymous):

\[f'(x)=ax^2-ax-2a\] and therefore \[f(x) = \frac{ax^3}{3}-\frac{ax^2}{2}-2ax+c\]

OpenStudy (anonymous):

good morning.

OpenStudy (amistre64):

howdy :)

OpenStudy (anonymous):

i assume this is looking for the cubic so the derivative is quadratic. with zeros at -1 and 2

OpenStudy (amistre64):

"please help to find cubic equation when f' (x) =0 at(-1;-3,5)" ??

OpenStudy (anonymous):

i am assuming you are looking for the cubic function whose derivative is zero at -1

OpenStudy (anonymous):

in other words \[f(x)=a_3x^3+a_2x^2+a_1x+a_0\]

OpenStudy (anonymous):

\[f(-1)=-3.5\] \[f(0)=0\] which of course means \[a_0=0\]

OpenStudy (amistre64):

a quartic f(x) can have 3 zeros if its a touch and turn, coupled with a thru and thru

OpenStudy (amistre64):

that would have 3 f' zeros then

OpenStudy (anonymous):

it is a cubic, and they have given us the zeros. the original function is a cubic. the derivative is quadratic. the zeros of the original function are 0, -2 and 4, which means it is \[f(x)=a(x+2)(x-4)x\]

OpenStudy (anonymous):

ho ho , that is a 4th degree poly. this one is sposed to be cubic!

OpenStudy (amistre64):

how does a cubic have 3 zeros in the derivative? unless the original question is in error....

OpenStudy (anonymous):

the derivative only has two zeros: -1 and 2

OpenStudy (amistre64):

"please help to find cubic equation when f' (x) =0 at(-1;-3,5" from the original statement ... I find this to either be an error, or the intended question...

OpenStudy (anonymous):

it is a parabola facing down, negative until -1, positive between -1 and 2, then negative

OpenStudy (amistre64):

f' = 0 at -1,-1,5 is my best interp of it.... right or wrong lol

OpenStudy (anonymous):

no it is not error. what it means is that f'(-1) = 0 and f(-1)=-3.5

OpenStudy (amistre64):

ooohhh....... that helps, thnx :)

OpenStudy (anonymous):

now let me see if i can finish this. before the day is over. lol

OpenStudy (anonymous):

ok where were we before we were interrupted?

OpenStudy (anonymous):

oh yes, you have a cubic polynomial with zeros at -2,0, and 4 so it must look like \[f(x)=a(x+2)(x-4)x\] where a is some negative number. negative because from the sign of the derivative we know that this polynomial is decreasing then increasing then decreasing. so we just need to find a

OpenStudy (anonymous):

The question says that f'(x) is 0 at -1,-3 and 5. But f'(-1)=0 yields a=0!

OpenStudy (anonymous):

i guess i cannot read. i read it as \[f'(-1)=0\] at \[(-1,-3.5)\]

OpenStudy (anonymous):

Yes but \[f^\prime(-1)=0 \implies a=0\]

OpenStudy (anonymous):

actually i get that too. damn

OpenStudy (anonymous):

the original quest. says find the equation of the cubic function when f'(x) = 0 at (-1;-3.5) then f'(x)<0 for x<-1 and x>2 f'(x)>0 for -1<x<2 f(0) =0 f(-2)=0 and f(4)=0

OpenStudy (anonymous):

in fact it get \[-6a=-8a\]

OpenStudy (anonymous):

@satellite73 that means a=0.

OpenStudy (anonymous):

think too many constraint to solve this one. yes, know it means a=0 that was what was bothering me.

OpenStudy (amistre64):

if a = 0, just use a different letter lol

OpenStudy (anonymous):

lets go slow. do we agree that \[f(x)=a(x+2)(x-4)x\] ?

OpenStudy (anonymous):

because f is cubic and the roots are given.

OpenStudy (anonymous):

The cubic is \[f(x)=ax(x+2)(x-4)\].

OpenStudy (anonymous):

ok so far so good. and we agree that a < 0 yes?

OpenStudy (anonymous):

thx a mil for the awesome insight, any chance i would be able to graph this equation.....

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

because of the behavior of the derivative: negative then positive then negative means f is decreasing then increasing then decreasing. so a < 0

OpenStudy (anonymous):

to show max an min values and intercepts

OpenStudy (anonymous):

and we also agree that \[f(x)=ax^3-2ax^2-8ax\]

OpenStudy (anonymous):

i know what they want you to graph, but i think this problem is crap. could be wrong. they want you to graph a cubic that is decreasing until -1, then increasing until 2, then decreasing. and they want it to cross the x axis at -2, 0, and 4.

OpenStudy (anonymous):

-4 that is

OpenStudy (anonymous):

no it says 4

OpenStudy (anonymous):

f(4)=0

OpenStudy (amistre64):

when f'(-1) = 0 and f(-3.5)=0 f'(x) = - ; (-inf,-1) U (2,inf) f'(x) = +; (-1,2) <.......-3.5..................-1.....0............2............> + 0 + 0 - - 0 + ????????

OpenStudy (amistre64):

little backwards, but still i quandry

OpenStudy (anonymous):

I think -1,-3.5 and the f'(x) being 0 date are wrong.

OpenStudy (anonymous):

it says f'(-1)=0 it changes direction at (-1,-3.5)

OpenStudy (amistre64):

f'(x)=0 could also indicate an inflection point... right?

OpenStudy (anonymous):

it is entirely possible that this thing is over determined. you have not only the zeros of the cubic, but also the zeros of its derivative. i will try it later because i gotta go, but i think there is a contradiction here.

OpenStudy (anonymous):

no it is not an inflection point because the derivative changes sign there.

OpenStudy (amistre64):

if the slope is the same on both sides of f' = 0, that indicates inflection i believe

OpenStudy (amistre64):

the signs dont change at -3.5

OpenStudy (anonymous):

derivative changes sign. that is what it says.

OpenStudy (anonymous):

is inflexion point noy determind by f''?

OpenStudy (amistre64):

in inflection point is made sure by f''; but an inflection point can still have a 0 slope in f'

OpenStudy (anonymous):

says f'(x)<0 for x<-1 f'(x)>0 for -1<x<2 f"(x)< 0 for x>2

OpenStudy (anonymous):

derivative changes sign so f changes direction from decreasing to increasing. still i cannot come up with a. will try later.

OpenStudy (amistre64):

"when f'(x) = 0 at (-1;-3.5) then f'(x)<0 for x<-1 and x>2 f'(x)>0 for -1<x<2"

OpenStudy (anonymous):

-3.5 is an inflection point..

OpenStudy (anonymous):

thx for all the help, egan helmie, south africa grade 12 -2011-provincial assesment task

OpenStudy (amistre64):

I see f(x) looking somthing like this if i interpret your info correctly

OpenStudy (anonymous):

oh no amistre, this is not a cubic polynomial. i interpret the question to be for a cubic polynomial. cubics do not look like this. also the second derivative of a cubic polynomial is a line, so only one inflection point, where the line crosses the x - axis (i.e. f''(x)=0). i am thinking this question was made up without sufficient thought. you cannot specify the zeros of a cubic polynomial, its value at some point, AND the zeros of the derivative.

OpenStudy (anonymous):

in fact i assert that there is no polynomial of degree 3 with zeros at -2, 0 and 4 for which f'(-1)=0. let me see if i fall flat on my face with a (rather simple) proof. by the factor theorem we know that \[f(x)=ax(x+2)(x-4)=ax^3-2ax^2-8ax\] thus \[f'(x)=3ax^2-4ax-8a\] and \[f'(-1)=3a+4a-8a\] the only way for this to be 0 is for \[a=0\] which is a contradiction.

OpenStudy (amistre64):

i trust your interpretations and its method; I dont trust that the given information is rigid enough to come to a solid conclusion tho ;)

OpenStudy (anonymous):

it is always possible i misinterpreted the question. but it is also possible some math teacher said "gee i will give them a problem to see if they can graph it" without thinking through whether such a thing exists, and that is the last second i am going to think about this.